# A Dipolar Chiral Spin Liquid on the Breathed Kagome Lattice

### 26 Mar 2026 v1

@@ Abstract
While this mechanism appears straightforward, quantum spin liquids (QSLs) often inhabit a small portion of a much broader and complex phase diagram [5]. This has motivated the investigation of a wide range of theoretical models, using beyond nearest neighbor interactions [7, 8], complicated multi-body terms [9, 10], and different lattices [11]. The identification of these models has laid out a blueprint for the search of quantum spin liquid candidate materials such as $\alpha$-RuCl3, Herbertsmithite and organic salts [12, 13, 14]. However, the lack of clear positive signatures has precluded the definite experimental identification of a spin liquid [15], motivating the search for new approaches and models for their stabilization and study.
Figure 1: (a) Schematic of a tweezer array platform where the position of the dipolar spins can be arbitrarily chosen. (b) Top view of the the two-dimensional breathed Kagome lattice, where the distance between the Kagome unit cells is scaled by the breathing parameter $\beta$. We label the small and large triangles by $\blacktriangle$ and $\bigtriangledown$, respectively. (c) A cluster composed of $N=75$ spins that exhibits signatures of the CSL: a lack of magnetic ordering (in single and two-site spin operators), time reversal symmetry breaking (i.e. non-zero chirality), and additional static and dynamic responses (see Figs. 9 and 10). (d) Phase diagram with respect to breathing $\beta$ and the strength $h_{z}$ of a local magnetic field pattern. There is a direct, second order transition between the paramagnet and the CSL at $h_{z}\sim 0.01$ for $\beta=1.5$ (yellow dashed line, details in Fig. 7.
@@
This work is organized as follows. In Sec. I, we briefly comment on the origin and properties of the CSL phase. In Sec. II, we introduce the model we consider: dipolar interacting spin-1/2 particles on a breathed Kagome lattice. The breathing parameter $\beta$ controls the scaling factor of the length of the basis vectors [Fig. 1(b)]. We also discuss the details of the numerical simulations. In Sec. III, we carefully characterize the presence of the CSL at a particular breathing parameter ($\beta=1.5$), highlighting the presence of time reversal symmetry (TRS) breaking and topological order. We then consider the broader breathing phase diagram in Sec. IV. This motivates the development of an effective low-energy microscopic description based on a perturbative analysis of the large breathing limit $\beta\to\infty$, which can quantitatively capture all our observations. In Sec. V, we study the extended phase diagram to probe the stability of the CSL and the feasibility of its adiabatic preparation in AMO systems. We find that the system naturally displays a 2nd order continuous phase transition between the CSL state and a paramagnet—opening the doors to its experimental preparation. Finally, in Sec. VI, we demonstrate that the stability of this phase extends to small finite-size clusters, which can be prepared in current experiments. Crucially, the direct access to the system’s local degrees of freedom enables the characterization of the TRS-breaking properties of this state as well as its topology; this manifests itself in robust chiral edge modes that can be directly accessed in experiments.
@@ I Background
While these models better capture real physical systems, they either require engineering complicated spin interactions (i and ii) or a spinful itinerant electron system (iii). In addition, the spin liquid phase only appears for a small sliver of the theoretical phase diagram, requiring precise control over the interactions. Recent progress towards frustrated Hubbard models in stacked 2D materials offer a tantalizing opportunity to observe the CSL [27], yet the realization of strong TRS-breaking fields or controllable next-neighbor interactions remains an important open challenge in both solid-state and AMO systems.
@@
II Theoretical Model
@@
When extending MPS methods to 2D, the linear structure of the MPS must be “wrapped” along the two-dimensional plane. Using this approach, the 2D model is directly mapped onto a longer-range 1D model at the cost of higher computational complexity. More precisely, when considering a cylinder geometry, the MPS’s bond dimension must increase exponentially with the width of the cylinder to preserve the accuracy of the calculation. This limits the calculations to cylinders of moderate width: in this work we focus on widths of 4, 5, and 6 unit cells. With the exception of Section VI, we use the infinite DMRG algorithm, which assumes the MPS to be translationally invariant along the cylinder axis, enabling a direct study of the cylinder’s thermodynamic limit.
@@
III Characterizing the Chiral Spin Liquid
@@
Indeed, for both $S_{x}$ and $S_{z}$, the onsite expectation value is zero [Fig. 2(a)], and it exhibits neither long-range nor algebraic order [Fig 2(b)]. We can complement this analysis by directly computing the MPS’s correlation length [39], observing that it does not increase with increasing bond dimension: this feature highlights the presence of a short-range correlated state as expected for a gapped phase of matter. In addition, the translation and rotation symmetries of the Kagome unit cell are also preserved; the pattern of two-point correlation functions (shading of the bonds) is symmetric across the entire lattice (up to small corrections of magnitude $\lesssim 2\times 10^{-2}$). Curiously, even though our model is highly anisotropic, the state’s spin-spin correlations exhibit a near SU(2) symmetry with $\expectationvalue{S^{x}_{i}S^{x}_{j}}$ and $\expectationvalue{S^{z}_{i}S^{z}_{j}}$ differing at most by $6.5\times 10^{-3}$. We return to this point later in our work when discussing our effective two-spin model, where the origin of this feature becomes more transparent [Sec. IV].
@@
While the previous considerations suggest that the ground state is a CSL, local correlations cannot alone unambiguously determine the presence of topological order [42]. Such confirmation requires the computation of non-local quantities. In the following sections we study four such probes: the fractionally quantized conductivity, the presence of anyonic flux as measured by the appropriate Wilson loops, the entanglement spectrum, and the modular matrices $\mathcal{S}$ and $\mathcal{U}$.
Conductivity and ground state degeneracy
The CSL is predicted to feature a quantized spin-Hall response, in which a gradient in a Zeeman field drives spin-current in the transverse direction. For numerical purposes, the spin-Hall response is best detected via Laughlin’s “flux-threading” argument [43]. By adiabatically threading spin-flux (via a twisted boundary condition), the amount of charge pumped through the system indicates its conductivity.
@@
Detecting anyonic flux through Wilson loops
We can corroborate the existence of topological sectors by computing the Wilson loop around the cylinder width. Building on previous work [23], we define the Wilson loop along a path $\ell=(a_{1},b_{1},a_{2},b_{2},\ldots)$ as the product of sequential raising and lowering operators:
$W_{\ell}=\prod_{i=1}^{|\ell|/2}S^{+}_{a_{i}}S^{-}_{b_{i}}\penalty 10000\ .$ (3)
where $W_{0}$ denotes the magnitude of the Wilson loop, and $\alpha$ encodes the phase which determines the flux information. Crucially, we observe that $|\langle W_{\ell}\rangle_{\text{ren}}|\gtrsim 0.3$ for all paths considered, suggesting the system is indeed in the deconfined regime. Moreover, the phase information of $\langle W_{\ell}\rangle_{\text{ren}}$ directly measures the anyonic flux in the system: computing it for the two wavefunctions at $\theta=0$ and $\theta=2\pi$, we get symmetric values, $\langle W_{\ell}\rangle_{\text{ren}}^{\theta=0}=0.81$ and $\langle W_{\ell}\rangle_{\text{ren}}^{\theta=2\pi}=-0.80$, corresponding to a flux of $\alpha^{\theta=0}=0$ and $\alpha^{\theta=2\pi}=\pi$, respectively. Note that while there is a phase ambiguity in the definition of $W_{\ell}$, this does not change the phase difference $\Delta\alpha$ between the two states which captures the braiding statistics.
Determining the nature of the chiral edge theory
One way of removing some of the sector ambiguity is by considering the entanglement spectrum of the system. Besides providing a direct identification of the ground state sector, it captures the spectrum of the edge mode [48, 49, 50]. This enables us to utilize our iMPS to directly characterize the edge theory that would occur in a finite sized boundary and extract the features of a gapless edge spectrum from a gapped bulk wavefunction.
@@
Characterizing anyonic statistics
Our analysis of the topological degeneracy reveals a fractionalized $S=1/2$ excitation but does not determine its braiding statistics. To this end, we compute the action of the modular transformations in the ground state manifold; these transformations directly connect the overlap of the ground states (under a suitable geometric transformation) to properties of the anyonic excitations [51, 54, 52]. The breathed Kagome lattice preserves the $\mathrm{C}_{3v}$ symmetry of the lattice, so the natural transformations to consider are $2\pi/3$ rotations, $\mathcal{R}_{2\pi/3}$, which can be cast as the product of the two generators of the modular transformations, $\mathcal{S}$ and $\mathcal{U}$ [39]:
@@
$US=e^{-i\frac{2\pi}{24}c}\begin{bmatrix}\theta_{1}S_{11}&\theta_{1}S_{1s}\\ \theta_{s}S_{s1}&\theta_{s}S_{ss}\end{bmatrix}=\frac{e^{-i\frac{2\pi}{24}(-1+0.37)}}{\sqrt{2}}\times$
$\matrixquantity[(1-0.006)&(1+0.004)e^{-i0.009}\\ (-i)(1+0.039)e^{i0.020}&(-i)(-1-0.002)e^{-i0.002}]$ (6)
@@
IV Breathing $\beta$ phase diagram
Figure 3: (a) Phase diagram with respect to the breathing parameter $\beta$ as measured by the (1) chirality $\langle\chi_{\blacktriangle}\rangle$, (2) translation symmetry breaking order parameter $O^{\circ/\diamondsuit_{SSB}}$, and (3) correlation length $\xi$. The nature of the transition is determined by the behavior of the order parameter $\chi_{\blacktriangle}$ and $\xi$. The transition from the KSL into the CSL at $\beta=1.3$ appears second order owing to the continuous behavior of the chirality and the associated diverging correlation length across the transition. These considerations also apply to the Glider-VBS to VBS transition at $\beta=2.5$. The transition from the CSL into the Glider-VBS pattern appears first order due to the discontinuous change of the chirality and the lack of a diverging correlation length. (b) Correlations for the different observed ground states at the $\beta$ values indicated by the colored vertical lines: (b1) the Kagome Spin Liquid (KSL), (b2) Glider-VBS, and (b3) VBS phases. The KSL exhibits homogeneous spin-spin correlations, without breaking any spin rotation or lattice translation. In the Glider-VBS phase, the correlations no longer obey translation symmetry, but instead obey a glider symmetry composed of a translation and a mirror transformation (dashed line). In the VBS state, this glider symmetry is also broken.
@@ IV Breathing $\beta$ phase diagram
As we approach the Kagome point $\beta\leq 1.3=\beta^{c}_{KSL-CSL}$, TRS is restored once again and the chirality operator no longer exhibits long-range order [Fig. 3(a)] Curiously, much like in the CSL state, the spin-spin correlations are featureless and exhibit no symmetry breaking pattern (in either spin or real space), suggesting the presence of a distinct spin liquid state [Fig. 3(b1)]. This observation is consistent with previous literature reporting a gapless Kagome spin liquid (KSL) in the Heisenberg antiferromagnet in the Kagome lattice [38], which is not destabilized by the long-range XY nature of our interactions [29].
Zooming in on the transition between the CSL and the KSL, our numerical exploration indicates the presence of a second-order phase transition. This is shown by two observations. First, around the transition $\beta\approx\beta^{c}_{\text{KSL-CSL}}$ we observe a divergence of the correlation length that is limited by the bond dimension of the iMPS wave function [Fig. 3(a)], consistent with a continuous phase transition. This also agrees with previous studies, which observed a second-order transition and whose understanding arises from the opening of the KSL spinon Dirac points [38, 29]. Second, the expectation value of the chirality smoothly interpolates from zero to its maximal value at $\beta\approx 1.5$.
On the other side of the CSL phase in the large breathing limit, we find evidence for two distinct VBS phases which preserve spin-rotation symmetry. When $2.1\lesssim\beta\lesssim 2.5$, we find that the system breaks translation symmetry while preserving a glide reflection [combination of translation and mirror transformation whose axis is the dashed line in Fig. 3(b2)]– we term this phase the Glider-VBS. We also observe that the chirality decays to zero with increasing bond dimension, suggesting that the true ground state in this regime respects TRS.
In the second phase for $\beta\gtrsim 2.5$, the system breaks both symmetries due to the unit cell doubling and all point group symmetries being broken. These two VBS phases are characterized by operators within a single triangle $|\langle O^{\circ}_{\blacktriangle}\rangle|$ and between two triangles $|\langle O^{\diamondsuit}_{\blacktriangle,\vec{r}}\rangle|$, as plotted in Fig. 3[(a)] (see SM for precise definition)[39].
@@
Understanding $\beta\to\infty$: building an effective spin-chirality model
The breathing parameter $\beta$ also provides a new perspective into the nature of the chiral spin liquid state. This perspective arises from the identification of the relevant degrees of freedom in the large breathing limit. For $\beta\to\infty$, the dominant energy scale $\approx(J/a^{3})$ arises from the interactions between spins in the same triangle, while the long-range inter-triangle interaction is weak $\approx(J/a^{3})\beta^{-3}$. In this limit, one can build an effective description of the system in terms of the low-energy subspace of each triangle. Surprisingly, this approach is able to capture the full phase diagram of the model, even at small $\beta$, suggesting its validity lies beyond a perturbative analysis around $\beta\to\infty$. At the same time, these effective degrees of freedom offer a new language with which to interpret the emergence of TRS-breaking, and shed light into the nature of the spin liquid state.

### 15 Apr 2026 v2

@@ Abstract
While this mechanism appears straightforward, quantum spin liquids (QSLs) often inhabit a small portion of a much broader and complex phase diagram [5]. This has motivated the investigation of a wide range of theoretical models, using beyond nearest neighbor interactions [7, 8], complicated multi-body terms [9, 10], and different lattices [11]. The identification of these models has laid out a blueprint for the search of quantum spin liquid candidate materials such as $\alpha$-RuCl3, Herbertsmithite, other Kagome materials, and organic salts [12, 13, 14, 15]. However, the lack of clear positive signatures has precluded the definite experimental identification of a spin liquid [16], motivating the search for new approaches and models for their stabilization and study.
Figure 1: (a) Schematic of a tweezer array platform where the position of the dipolar spins can be arbitrarily chosen. (b) Top view of the the two-dimensional breathed Kagome lattice, where the distance between the Kagome unit cells is scaled by the breathing parameter $\beta$. We label the small and large triangles by $\blacktriangle$ and $\bigtriangledown$, respectively. (c) A cluster composed of $N=75$ spins that exhibits signatures of the chiral spin liquid (CSL): a lack of magnetic ordering (in single and two-site spin operators), time reversal symmetry breaking (i.e. non-zero chirality), and additional static and dynamic responses (see Figs. 9 and 10). (d) Phase diagram with respect to breathing $\beta$ and the strength $h_{z}$ of a local magnetic field pattern. There is a direct, second order transition between the paramagnet and the CSL at $h_{z}\sim 0.008$ for $\beta=1.5$ (yellow dashed line, details in Fig. 8).
@@
After a brief introduction to the CSL phase in Sec. I, our work is organized across two sections. In Sec. II, we introduce the central model in our work—dipolar interacting spin-1/2 particles on a breathed Kagome lattice—and characterize the resulting CSL state. We present a careful study of the CSL state (both TRS symmetry breaking and topological order), the phase diagram of our model as a function of breathing, and introduce an effective model utilizing the low-energy degrees of freedom within each unit-cell.
In Sec. III, we discuss how the tools of AMO platforms enable us to better stabilize, prepare, and study the CSL in experiments. We begin by studying the CSL state under different finite-sized geometries and using additional Ising interactions, highlighting how geometry control and Floquet engineering can enhance the stability of the CSL phase. We then investigate how additional fields can be utilized to engineer an adiabatic path into the CSL state. While global fields lead to a complex ladder of magnetically ordered phases, a local pattern enables a single, direct transition between a trivial paramagnetic and the CSL state. Finally, we study how direct access to the system’s local degrees of freedom enables the characterization of CSL state through a transverse spin response, a robust chiral edge mode, and signatures of the trapping of a semion. We conclude by remarking on how our protocol can be implemented in state-of-the-art experimental platforms in Rydberg atom and molecular tweezer arrays, as well as neutral atoms in optical lattices.
@@ I Background
While these models better capture real physical systems, they either require engineering complicated spin interactions (i and ii) or a spinful itinerant electron system (iii). In addition, the spin liquid phase only appears for a small sliver of the theoretical phase diagram, requiring precise control over the interactions. Recent progress towards frustrated Hubbard models in stacked 2D materials offer a promising route to observe the CSL [28, 29], yet the realization of strong TRS-breaking fields or controllable next-neighbor interactions remains an important open challenge in both solid-state and AMO systems.
@@
II Dipolar Spins in a breathed Kagome Lattice
Theoretical Model
@@ II Dipolar Spins in a breathed Kagome Lattice
When extending MPS methods to 2D, the linear structure of the MPS must be “wrapped” along the two-dimensional plane. Using this approach, the 2D model is directly mapped onto a longer-range 1D model at the cost of higher computational complexity. This limits the calculations to cylinders of moderate width: in this work we focus on widths of 4, 5, and 6 unit cells. Unless when discussing finite clusters, we use the infinite DMRG algorithm, which assumes the MPS to be translationally invariant along the cylinder axis, enabling a direct study of the cylinder’s thermodynamic limit.
@@
II.1 Characterizing the Chiral Spin Liquid
@@
Indeed, for both $S_{x}$ and $S_{z}$, the onsite expectation value is zero [Fig. 2(a)], and it exhibits neither long-range nor algebraic order [Fig 2(b)]. We can complement this analysis by directly computing the MPS’s correlation length [41], observing that it does not increase with increasing bond dimension: this feature highlights the presence of a short-range correlated state as expected for a gapped phase of matter. In addition, the translation and rotation symmetries of the Kagome unit cell are also preserved; the pattern of two-point correlation functions (shading of the bonds) is symmetric across the entire lattice (up to small corrections of magnitude $\lesssim 2\times 10^{-2}$). Curiously, even though our model is highly anisotropic, the state’s spin-spin correlations exhibit a near SU(2) symmetry with $\expectationvalue{S^{x}_{i}S^{x}_{j}}$ and $\expectationvalue{S^{z}_{i}S^{z}_{j}}$ differing at most by $6.5\times 10^{-3}$. We return to this point later in our work when discussing our effective two-spin model, where the origin of this feature becomes more transparent [Sec. II.2].
@@
While the previous considerations suggest that the ground state is a CSL, local correlations cannot alone unambiguously determine the presence of topological order [44]. Such confirmation requires the computation of non-local quantities. In the following sections we study four such probes: the fractionally quantized conductivity, the entanglement spectrum, and the modular matrices $\mathcal{S}$ and $\mathcal{U}$.
Conductivity and ground state degeneracy—The CSL is predicted to feature a quantized spin-Hall response, in which a gradient in a Zeeman field drives spin-current in the transverse direction. For numerical purposes, the spin-Hall response is best detected via Laughlin’s “flux-threading” argument [45]. By adiabatically threading spin-flux (via a twisted boundary condition), the amount of charge pumped through the system indicates its conductivity.
@@
Determining the nature of the chiral edge theory—One way of removing some of the sector ambiguity is by considering the entanglement spectrum of the system. Besides providing a direct identification of the ground state sector, it captures the spectrum of the edge mode [49, 50, 51]. This enables us to utilize our iMPS to directly characterize the edge theory that would occur in a finite-sized boundary and extract the features of a gapless edge spectrum from a gapped bulk wavefunction.
@@
Characterizing anyonic statistics—To this end, we compute the action of the modular transformations in the ground state manifold; these transformations directly connect the overlap of the ground states (under a suitable geometric transformation) to properties of the anyonic excitations [52, 55, 53]. The breathed Kagome lattice preserves the $\mathrm{C}_{3v}$ symmetry of the lattice, so the natural transformations to consider are $2\pi/3$ rotations, $\mathcal{R}_{2\pi/3}$, which can be cast as the product of the two generators of the modular transformations, $\mathcal{S}$ and $\mathcal{U}$ [41]:
@@
$US=e^{-i\frac{2\pi}{24}c}\begin{bmatrix}\theta_{1}S_{11}&\theta_{1}S_{1s}\\ \theta_{s}S_{s1}&\theta_{s}S_{ss}\end{bmatrix}$= (5)
$\frac{e^{-i\frac{2\pi}{24}(-1{\scriptscriptstyle+0.37})}}{\sqrt{2}}\begin{bmatrix}1{\scriptstyle-0.006}&(1{\scriptstyle+0.004}){\scriptstyle e^{-i0.009}}\\ -i(1{\scriptstyle+0.039}){\scriptstyle e^{i0.020}}&-i(-1{\scriptstyle-0.002}){\scriptstyle e^{-i0.002}}\end{bmatrix}$
@@
Figure 3: (a) Phase diagram with respect to the breathing parameter $\beta$ as measured by the (1) chirality $\langle\chi_{\blacktriangle}\rangle$, (2) rotation and glider symmetry breaking order parameter $O^{\text{rot}}$ and $O^{\text{glid}}$, and (3) correlation length $\xi$. The nature of the transition is determined by the behavior of the order parameter $\chi_{\blacktriangle}$ and $\xi$. The transition from the KSL into the CSL at $\beta=1.3$ appears second order owing to the continuous behavior of the chirality and the associated diverging correlation length across the transition. These considerations also apply to the valence-bond solid (VBS) phases at the Glider-VBS to VBS transition at $\beta=2.5$. The transition from the CSL into the Glider-VBS pattern appears first order due to the discontinuous change of the chirality and the lack of a diverging correlation length. (b) Correlations for the different observed ground states at the $\beta$ values indicated by the colored vertical lines: (b) the Kagome Spin Liquid (KSL), (c) Glider-VBS, and (d) VBS phases. The KSL exhibits homogeneous spin-spin correlations, without breaking any spin rotation or lattice translation. In the Glider-VBS phase, the correlations no longer obey translation symmetry, but instead obey a glider symmetry composed of a translation and a mirror transformation (dashed line). In the VBS state, this glider symmetry is also broken.
@@
II.2 Breathing $\beta$ phase diagram
@@
As we approach the Kagome point $\beta\leq 1.3=\beta^{c}_{KSL-CSL}$, TRS is restored once again and the chirality operator no longer exhibits long-range order [Fig. 3(a)] Curiously, much like in the CSL state, the spin-spin correlations are featureless and exhibit no symmetry breaking pattern (in either spin or real space), suggesting the presence of a distinct spin liquid state [Fig. 3(b)]. This observation is consistent with previous literature reporting a gapless Kagome spin liquid (KSL) in the Heisenberg antiferromagnet in the Kagome lattice [40], which is not destabilized by the long-range XY nature of our interactions [31].
Zooming in on the transition between the CSL and the KSL, our numerical exploration indicates the presence of a second-order phase transition. This is shown by two observations. First, around the transition $\beta\approx\beta^{c}_{\text{KSL-CSL}}$ we observe a divergence of the correlation length that is limited by the bond dimension of the iMPS wave function [Fig. 3(a)], consistent with a continuous phase transition. This also agrees with previous studies, which observed a second-order transition and whose understanding arises from the opening of the KSL spinon Dirac points [40, 31]. Second, the expectation value of the chirality smoothly interpolates from zero to its maximal value by $\beta\approx 1.5$.
On the other side of the CSL phase in the large breathing limit, we find evidence for two distinct valence-bond solid (VBS) phases which preserve spin-rotation symmetry. When $2.1\lesssim\beta\lesssim 2.5$, we find that the system breaks rotation and translation symmetries—captured by the non-zero value of $O^{\text{rot}}$ (see SM for details [41])—while preserving a glide transformation [combination of translation and mirror transformation whose axis is the dashed line in Fig. 3(c)]– we term this phase the Glider-VBS. We also observe that the chirality decays to zero with increasing bond dimension, suggesting that the true ground state in this regime respects TRS.
In the second phase for $\beta\gtrsim 2.5$, the gliding symmetry is also broken. We introduce $O^{\text{glid}}$ (see SM for details [41])) as an order parameter to distinguish these two phases, and observe it taking a non-zero value only in the VBS phase.
@@
II.3 Building an effective spin-chirality model
The breathing parameter $\beta$ also provides a new perspective into the nature of the chiral spin liquid state. This perspective arises from the identification of the relevant degrees of freedom in the large breathing limit. For $\beta\to\infty$, the dominant energy scale $\approx(J/a^{3})$ arises from the interactions between spins in the same triangle, while the long-range inter-triangle interaction is weak $\approx(J/a^{3})\beta^{-3}$. In this limit, one can build an effective description of the system in terms of the low-energy subspace of each triangle. Surprisingly, this approach is able to capture the full phase diagram of the model, even at small $\beta$, suggesting its validity lies beyond a perturbative analysis around $\beta\to\infty$. At the same time, these effective degrees of freedom offer a new language with which to interpret the emergence of TRS-breaking, and offer a path towards understanding the mechanisms generate the spin liquid state.